# N-DriverMotion: Driver motion learning and prediction using an event-based camera and directly trained spiking neural networks

### v1

@@ Abstract
Driver motion recognition is a principal factor in ensuring the safety of driving systems. This paper presents a novel system for learning and predicting driver motions and an event-based high-resolution (1280x720) dataset, N-DriverMotion, newly collected to train on a neuromorphic vision system. The system comprises an event-based camera that generates the first high-resolution driver motion dataset representing spike inputs and efficient spiking neural networks (SNNs) that are effective in training and predicting the driver’s gestures. The event dataset consists of 13 driver motion categories classified by direction (front, side), illumination (bright, moderate, dark), and participant. A novel simplified four-layer convolutional spiking neural network (CSNN) that we proposed was directly trained using the high-resolution dataset without any time-consuming preprocessing. This enables efficient adaptation to on-device SNNs for real-time inference on high-resolution event-based streams. Compared with recent gesture recognition systems adopting neural networks for vision processing, the proposed neuromorphic vision system achieves comparable accuracy, 94.04%, in recognizing driver motions with the CSNN architecture. Our proposed CSNN and the dataset can be used to develop safer and more efficient driver monitoring systems for autonomous vehicles or edge devices requiring an efficient neural network architecture.
@@ 1 Introduction
With the advancement of artificial intelligence (AI), it is being applied in various industrial fields, and among them, vehicle AI systems are emerging as one of the most prominent applications. AI used in vehicles is used as a system that assists autonomous driving or the safety of drivers and pedestrians by integrating it with the control system [2]. In particular, the EU and the United States have introduced mandatory requirements for various driver assistance systems to ensure safe road traffic by regulations on general automotive safety [1]. This has led to an urgent need to research and develop driver assistance systems (ADAS) using AI. In response to these demands, we propose N-DriverMotion, an event-based dataset for neuromorphic learning and predicting driver motions on an efficient convolutional spiking neural network (CSNN). This driver motion recognition research, comprising a neuromorphic framework for efficient CSNN configuration, implements driver safety assistance by incorporating a high-resolution event-based camera. The contributions of the proposed driver motion recognition system include the following:
@@
  • We create an event-based high-resolution (1280x720) dataset, N-DriverMotion, for large-scale driver motion recognition using an event-based camera: We believe that this paper is the first to study driver motion recognition using a high-resolution event camera and spiking neural networks. The event-based dataset presents 13 driver motion categories classified by direction (front, side), lighting (bright, moderate, dark), and participant.
  • We present a novel simplified four-layer convolutional spiking neural network (CSNN), directly trained by the high-resolution dataset without any time-consuming preprocessing. This enables efficient adaptation to on-device SNNs for real-time inference over high-resolution event-based camera streams. Furthermore, the proposed neuromorphic vision system achieves comparable accuracy, 94.04%, in recognizing driver motions with the CSNN architecture, developing safer and more efficient driver monitoring systems for autonomous vehicles or edge devices requiring an efficient neural network architecture.
  • We configured the efficient CSNN for driver motion recognition using one of the widely used practical neuromorphic frameworks, the Intel Lava neuromorphic framework  [^1] Lava is an open-source software framework for developing neuromorphic systems. Please refer to https://github.com/lava-nc/lava for real-world applications and adoptions with optimization of hyperparameter values.
@@
The remainder of this paper is organized as follows. Section 2 briefly reviews related work, while Section 3 describes the proposed driver motion recognition system. Section 4 presents the implementation results and performance evaluations and Section 5 concludes this paper.
@@ 2 RELATED WORK
Event-based cameras have advantages over frame-based cameras on properties such as high dynamic range, high temporal resolution, low latency, and low power consumption because of their sparsity in stream events when intensity changes [28]. Therefore, they are gradually used in machine-learning vision systems for gesture learning and recognition. Hand-gesture recognition [12], gesture and facial expression recognition [8], DECOLLE [7] for deep continuous local learning, EDDD for event-based drowsiness driving detection [14], TORE [6] for time-ordered recent events, event-based asynchronous sparse convolutional networks [5], and EST [4] for end-to-end learning of representations for asynchronous event-based data have been devised on deep learning-based neural networks (DNNs) and event representations, exploiting the sparsity and temporal dispersion of event-based gestures. Riccardo et al. trained a deep neural network with a pre-processed DVS hand gesture dataset for recognizing hand gestures on a Loihi neuromorphic processor [15]. The trained DNN was converted into the spike domain for deployment on Intel Loihi [29] for real-time gesture recognition.
@@
Unlike DNN-based approaches that require converting event-based datasets to static patterns for processing, recent research has shifted to direct training and learning on SNNs using event-based datasets. This is because events are inherently and naturally suited for SNNs operating in continuous time. Moreover, SNNs directly exploit the temporal information of events for energy-efficient computation. SLAYER [16] has proposed an error backpropagation mechanism for offline training on SNNs. It receives events and leverages the backpropagation method to train synaptic weights and axonal delays directly. Amir et al. proposed an end-to-end event-based gesture recognition system using an event-based camera and a TrueNorth processor configured with a convolutional neural network (CNN) [19]. It recognized hand gestures in real-time from hand gesture event streams captured by a Dynamic Vision Sensor(DVS).
@@
Event-based vision systems have been widely used in various applications, including vision systems for autonomous driving [11, 13] and object detection [9, 10]. These applications exploit event-based cameras to capture the spatial and temporal driving and object event data.
@@ 3 PROPOSED EVENT-BASED DRIVER MOTION RECOGNITION SYSTEM
Our proposed model is developed based on existing models [19, 16] for DVS Gesture recognition which implements SNN training and 3D spike convolution operation. The SLAYER and TrueNorth models consist of 8-layer and 16-layer, respectively. Figure 2 presents our proposed network model and system, where the network is simplified and comprised of only 4 layers. The 1280x720 resolution event streams for driver motion recognition were recorded by an event-based camera for 3 seconds, and delivered to the first 3D pooling layer. Unlike conventional convolutional neural networks, our model first implements a pooling layer before passing the data to the convolution layer. The main reason for this is to down-sample feature maps resolving a large memory usage, as well as minimizing the loss of data. The pooling layer uses a kernel of size 8*8 and passes to the convolution layer which uses a 5*5 kernel to produce 16 layers of features. The features are then fed to 2 final fully-connected layers which would provide 13 outputs representing the spike rates for the given event data.
@@ 4.1 Event-based camera and data conversion
The Prophesee’s Metavision EVK4222EVK4 camera is an HD Event-Based Vision evaluation kit mounted with IMX636 sensor, developed by Sony and Prophesee. Please refer to https://www.prophesee.ai/event-camera-evk4/ camera is one of the latest event-based vision sensors that supports up to 1280x720 pixel resolution. Generally, when there exists a change in pixel values exceeding a certain threshold defined by the user, the event-based camera detects such changes in brightness asynchronously and generates events specific to that pixel [28]. Each event contains pixel information relative to the position described in the x and y coordinates, indicating a motion’s change and a timestamp for the occurrence of the event. The device offers a high dynamic range (86 dB, however, it can reach over 120 dB based on low light cutoff measurement being: 0.08 lux) and a typical and maximum event rate of 1.06 giga-events per second (Geps). To allow direct usage of the spike data earned from the high-resolution event-based camera without requiring additional resources, we have minimized modification of the extractor such that it can retrieve coordinate, polarity, origin (newly implemented feature in EVK4), and timestamp information. Moreover, the most meaningful motion information was contained in the central 720x720 pixels so that the size of the input events was cropped to the 720x720 map size.
@@
4.2 Datasets
Over the past years, a profusion of gesture datasets captured with a simple frame-based sensor has been presented. However, to stimulate the improvement of event-based computer vision, Hu et al. [17] strongly support the importance of the DVS dataset. Serrano et al. introduced the first labeled event-based neuromorphic vision sensor dataset which originated from the classical MNIST digit recognition dataset by moving the images along certain directions within the screen [18], with being further developed by Orchard et al. where several frame artifacts were removed with a pan-tilt unit [22]. It could be observed that generating artificial movement to static images enabled the production of event outputs, which became a crucial factor for object recognition in neuromorphic systems such as spiking neural networks [23].
However, it is observed that static image recognition with event-based vision sensors is ineffective as the purpose of the usage is mainly focused on dynamic scenes. Recognizing such drawbacks, some datasets which consist of dynamic scenes have been presented. Berner et al. presented novel datasets that converted the existing visual video benchmarks for object tracking, and action/object recognition into spiking neuromorphic datasets with DAVIS camera’s output [20].
Even though these datasets show dynamic movements in a spiking neuromorphic way, according to Tan et al., a DVS camera produces microsecond temporal resolution output. In contrast, datasets produced by conventional vision tools such as video recorders or color cameras have tens of milliseconds which would cause loss in high temporal frequency during conversion. Moreover, they add that there is a high chance of subsidiary unwanted artifacts being contained during conversion [21].
To address such problems, Amir et al. present the DvsGesture dataset, which directly took the scenery with the DVS128 camera. Also, the DvsGesture datasets are taken at different light conditions as DVS sensors do not affect much from brightness and add some meaningful noises [19]. However, the DvsGesture dataset is mainly utilized for simple classification tasks, which could be considered less practical.
Number Gesture
@@ 12 Bends down (a: left, b: right)
Table 1: Classification of driver’s gestures.
@@
The N-DriverMotion dataset comprises 1,239 instances of a set of 13 driver motions as shown in Figure 3 and Table 1, which are composed of head falling (front, sides, and back), holding a cellphone(right/left hand), looking back(to the right/left side), the body going down(to the right/left side), and all the same motions above in side-captured manner. The dataset was collected from 23 subjects under 3 different lighting conditions related to brightness. Each subject was to perform all 13 motions in a single recording, where each motion was recorded around 3 seconds, under the same brightness condition. 3 brightness conditions were classified as "bright", "moderate", and "dark", as to test how much the brightness would affect the recordings. To evaluate the classifier’s performance, we randomly selected 992 motions for the training set, and the leftover 247 motions were designated as the test set.

### v2

@@ Abstract
Driver motion recognition is a key factor in ensuring the safety of driving systems. This paper presents a novel system for learning and predicting driver motions, along with an event-based (720x720) dataset, N-DriverMotion, newly collected to train a neuromorphic vision system. The system includes an event-based camera that generates a driver motion dataset representing spike inputs and efficient spiking neural networks (SNNs) that are effective in training and predicting the driver’s gestures. The event dataset consists of 13 driver motion categories classified by direction (front, side), illumination (bright, moderate, dark), and participant. A novel optimized four-layer convolutional spiking neural network (CSNN) was trained directly without any time-consuming preprocessing. This enables efficient adaptation to energy- and resource-constrained on-device SNNs for real-time inference on high-resolution event-based streams. Compared to recent gesture recognition systems adopting neural networks for vision processing, the proposed neuromorphic vision system achieves competitive accuracy of 94.04% in a 13-class classification task, and 97.24% in an unexpected abnormal driver motion classification task with the CSNN architecture. Additionally, when deployed to Intel Loihi 2 neuromorphic chips, the energy-delay product (EDP) of the model achieved 20,721 times more efficient than that of a non-edge GPU, and 541 times more efficient than edge-purpose GPU. Our proposed CSNN and the dataset can be used to develop safer and more efficient driver-monitoring systems for autonomous vehicles or edge devices requiring an efficient neural network architecture.
@@ 1 Introduction
With the advancement of artificial intelligence (AI), it is being applied across various industrial fields, and among these, vehicle AI systems are emerging as one of the most prominent applications. Onboard AIs in vehicles are used to assist autonomous driving and the safety of drivers and pedestrians by integrating with the control system [1]. In particular, the EU and the United States have introduced mandatory requirements for various driver assistance systems to ensure safe road traffic by regulations on general automotive safety [2]. The primary objective of this regulation is to enhance the protection of elderly drivers, vehicle occupants, pedestrians, and cyclists. Given that research shows human error is the cause of 95% of accidents, implementing this regulation is projected to save more than 25,000 lives and prevent at least 140,000 serious injuries by 2038. This has led to an urgent need to research and develop driver assistance systems (ADAS) using AI.
Conventional AI vision systems for autonomous vehicles utilize AI platforms for training with large datasets, after which the learned information is transferred to individual vehicles. These systems are typically implemented using GPU or AI accelerator-based hardware for data training and inference, which result in high power consumption, slow response times, and challenges in real-time prediction [3]. The exponential increase in neural network computations required for massive training datasets in autonomous driving has exacerbated these limitations. To address these challenges, a new generation of low-power, high-performance, and highly efficient AI vision systems is expected to become central to AI implementation in autonomous driving.
Neuromorphic AI vision systems, in particular, represent a way to support AI technology for autonomous vehicles that is gaining increasing significance [4]. Event cameras, also called neuromorphic cameras, and neuromorphic processors, which mimic the functionality of the human eye and brain, offer ultra-low power consumption, high efficiency, and rapid responsiveness. Consequently, these technologies are actively applied in various applications in autonomous vehicles, where they serve as core components for the sensory and processing needs [5, 6, 7].
In response to these demands, we propose N-DriverMotion, an event-based dataset and vision system for neuromorphic learning and predicting driver motions on an efficient convolutional spiking neural network (CSNN). This driver motion recognition research, comprising a neuromorphic framework for efficient CSNN configuration in terms of energy and latency, implements driver safety assistance by incorporating a high-resolution event-based camera. The contributions of the proposed driver motion recognition system include the following:
@@
  • We create an event-based dataset, N-DriverMotion (720 * 720 resolution), for large-scale driver motion recognition using an event-based camera: To the best of our knowledge, this is the first study for driver motion recognition using a high-resolution event camera and spiking neural networks. The event-based dataset presents 13 driver motion categories classified by direction (front, side), illumination (bright, moderate, dark), and participant.
[…]
  • We design an efficient CSNN for driver motion recognition for one of the widely used practical neuromorphic frameworks, the Intel Lava neuromorphic framework. We also propose a version of our CSNN for the Loihi 2 processor [8, 9]. We observed that the energy-delay product (EDP) of the proposed model on Loihi 2 exhibited 20,721 times more efficient than that of a non-edge GPU and 541 times more efficient than that of an edge purpose GPU, which is considered to be one of the most important aspects of edge-device AI.
We designed an optimized convolutional spiking neural network (CSNN) for efficient energy use and deployment on on-device applications and systems like Loihi 2. To train the proposed CSNN, we used a surrogate gradient descent-based backpropagation method, SLAYER [10], for direct training with our N-DriverMotion dataset. Additionally, instead of resizing the high-resolution event frames directly, we added a pooling layer to enable adaptive deployment on GPUs and the Loihi 2 system, depending on memory requirements.
@@
The remainder of this paper is organized as follows: Section 2 briefly reviews related work, while Section 3 describes the proposed driver motion recognition system. Section 4 presents the implementation results and performance evaluations and Section 4.7 concludes this paper.
@@ 2 RELATED WORK
Event-based cameras have advantages over frame-based cameras in properties such as high dynamic range, high temporal resolution, low latency, and low power consumption due to their sparse event streams when intensity changes [11]. Therefore, they have been increasingly adopted in machine learning vision systems for gesture recognition and learning. Hand-gesture recognition [12], gesture and facial expression recognition [13], DECOLLE [14] for deep continuous local learning, EDDD for event-based drowsiness driving detection [15], TORE [16] for time-ordered recent events, event-based asynchronous sparse convolutional networks [17], and EST [18] for end-to-end learning of representations for asynchronous event-based data have been developed on deep learning-based neural networks (DNNs) and event representations, exploiting the sparsity and temporal dispersion of event-based gestures. Riccardo et al. [6] trained a deep neural network on a preprocessed DVS hand gesture dataset for recognizing hand gestures on a Loihi neuromorphic processor. The trained DNN was converted into the spike domain for deployment on Intel Loihi [19] for real-time gesture recognition.
@@
Unlike DNN-based approaches, which require converting event-based datasets into static patterns for processing, recent research has shifted toward direct training and learning on SNNs using event-based datasets. This is because events are inherently suited for SNNs operating in continuous time. Moreover, SNNs directly exploit the temporal information of events for energy-efficient computation. SLAYER proposes an error backpropagation mechanism for offline training of SNNs. It receives events and leverages the backpropagation method to train synaptic weights and axonal delays directly. Amir et al. [20] proposed an end-to-end event-based gesture recognition system using an event-based camera and a TrueNorth processor configured with a convolutional neural network (CNN). It recognized hand gestures in real time from gesture event streams captured by a Dynamic Vision Sensor (DVS).
@@
Event-based vision systems have been widely used in various applications, including vision systems for autonomous driving [21, 22] and object detection [23, 24]. These applications exploit event-based cameras to capture the spatial and temporal event data of driving and objects.
Figure 1: Simplified structure of Loihi which illustrates how neurocore is structured and operates.
Loihi 2 is the second generation of Intel neuromorphic chip designed for energy-efficient AI applications [9]. Loihi chip applies Current Based Leaky-Integrate and Fire neuron (CUBA LIF), as its main neuron model. The overall behavior of CUBA LIF model is close to that of LIF, but is more biologically plausible as CUBA LIF model includes the temporal dynamics of the postsynaptic input current [25]. A single Loihi chip consists of 128 neurocores, which are groups of neurons that function according to the compartments. Each neurocore is able to communicate with other neurocores through the asynchronous Network-on-Chip (NoC) where the messages are packetized for sending and receiving. Although a single chip provides only 128 neurocores, Loihi is capable of transporting information without increasing latency for message exchange between multiple chips. The simplified model of Loihi is shown in Figure 1, which illustrates the asynchronous neuron model mechanism, along with the implementation of online learning or a continual learning mechanism. The Loihi chip has been used with the Lava software framework to support the development of agile neuromorphic software and applications [27, 26, 28, 29, 30, 31].
While event-based systems have shown promise in previous research for autonomous driving and gesture recognition, they still face challenges in achieving real-time performance with low power consumption in highly dynamic environments. Additionally, existing systems may struggle to handle the diverse and unpredictable nature of recognition scenarios, particularly when dealing with rapid motions or changing lighting conditions, which can generate massive event streams. Our proposed system with N-DriverMotion addresses these challenges by leveraging the energy-efficient CSNN architecture on Loihi 2. This enables fast, real-time processing of large-scale event data while maintaining low power consumption. Furthermore, the system ensures that event data from event cameras is processed promptly without bottlenecks, which is important for safety-critical applications like autonomous driving.
@@ 3 PROPOSED EVENT-BASED DRIVER MOTION RECOGNITION SYSTEM
Our proposed model is developed based on existing models [20, 10] for DVS Gesture recognition, which implements SNN training and a 3D spike convolution operation. The SLAYER and TrueNorth models consist of 8 layers and 16 layers, respectively. Figure 3 presents our proposed network model and system, where the network is simplified and comprises only 4 layers. When tested with a more complex model, we observed that it not only occupied more memory, but also had a detrimental impact on overall performance, as more layers hindered the neurons from firing spikes, leading to lower event rate because spikes could not be passed to consecutive neuron layers.
The 720x720 resolution event streams for driver motion recognition were recorded by an event-based camera for 3 seconds and delivered to the first 3D pooling layer. Unlike conventional convolutional neural networks, where convolution layers are processed first, our model implements a pooling layer before passing the data to the convolution layer. The main reason for this is to down-sample feature maps to resolve large memory usage, as well as to extract features. Compared to classical pooling layers, SNN pooling layers are considered trainable layer, as they not only down-sample the event streams but also take into account temporal aspects, enabling the learning of spatio-temporal features from the input. The pooling layer uses a kernel of size 8x8 and passes data to the convolution layer, which uses a 5x5 kernel to produce 16 channels of features. The features are then fed to 2 final fully-connected layers, which provide 13 outputs representing the spike rates for the given event data.
We notice that the existing spiking neural networks designed for other applications are not optimized for autonomous driving, which requires low power consumption, real-time inference, and efficient deployment on actual neuromorphic chips. To effectively implement driver motion recognition for autonomous driving, it is important to have a neural network that is optimized for neuromorphic hardware (simple but without performance degradation) as well as an event-based dataset for training such networks. We addressed these elements by designing, implementing, and testing a network that takes these features into consideration.
@@ 4.1 EVENT-BASED CAMERA AND DATA CONVERSION
Prophesee’s Metavision EVK4111EVK4 camera is an HD Event-Based Vision evaluation kit mounted with IMX636 sensor, developed by Sony and Prophesee. Please refer to https://www.prophesee.ai/event-camera-evk4/ camera is one of the latest event-based vision sensors that supports up to 1280x720 pixel resolution. Generally, when a change in pixel values exceeding a certain threshold defined by the user occurs, the event-based camera asynchronously detects such changes in brightness and generates events specific to that pixel [11]. Each event contains pixel information describing the position in the x and y coordinates, indicating a motion’s change, and a timestamp for the occurrence of the event. The device offers a high dynamic range (86 dB, however, it can reach over 120 dB based on low light cutoff measurement being: 0.08 lux) and a typical and maximum event rate of 1.06 giga-events per second (Geps).
In the experiment, in order to allow direct usage of the spike data obtained from the high-resolution event-based camera without requiring additional resources, we minimized modification to the extractor so that it can retrieve coordinate, polarity, origin (a newly implemented feature in the EVK4), and timestamp information. Moreover, we observed that the most meaningful motion information was contained in the central 720x720 pixels so that the size of the input events was cropped to 720x720 pixels.
4.2 DATASET
Over the past years, a profusion of gesture datasets captured with simple frame-based sensors has been presented. However, to stimulate the improvement of event-based computer vision, Hu et al. [37] strongly support the importance of the DVS dataset. Serrano et al. [38] introduced the first labeled event-based neuromorphic vision sensor dataset which originated from the classical MNIST digit recognition dataset by moving the images along certain directions within the screen. This dataset was further developed by Orchard et al. [39], who removed several frame artifacts using a pan-tilt unit. It can be observed that generating artificial movement for static images enabled the production of event outputs, which became important for object recognition in neuromorphic systems such as spiking neural networks [7]. However, it has been observed that static image recognition with event-based vision sensors is ineffective, as their main purpose is focused on dynamic scenes. Recognizing these drawbacks, some datasets consisting of dynamic scenes have been presented. Berner et al. [40] introduced novel datasets that converted the existing visual video benchmarks for object tracking and action/object recognition into spiking neuromorphic datasets using the DAVIS camera’s output.
@@ 4.2 DATASET
Table 1: Labels of the corresponding driver gestures
Label Gesture
@@ 12 Bends down (a: left, b: right)
Even though these datasets show dynamic movements in a spiking neuromorphic way, according to Tan et al. [41], a DVS camera produces microsecond temporal resolution output. In contrast, datasets produced by conventional vision tools such as video recorders or color cameras have tens of milliseconds of temporal resolutions, which causes a loss of high temporal frequency during conversion. Moreover, they add that there is a high chance of subsidiary unwanted artifacts being present during conversion. To address these problems, Amir et al. [20] present the DvsGesture dataset, which directly captured scenes with the DVS128 camera. Additionally, the DvsGesture datasets were captured under different lighting conditions, as DVS sensors are less affected by brightness and introduce some meaningful noise. However, the DvsGesture dataset is mainly utilized for simple classification tasks, which may be considered less practical.
@@
The N-DriverMotion dataset comprises 1,239 instances of a set of 13 driver motions (1 normal driving motion and 12 driving motions that are considered possibly dangerous) as shown in Figure 4 and Table 1. These motions include head falling (front, sides, and back), holding a cellphone (right/left hand), looking back (to the right/left side), the body going down (to the right/left side), and all the same motions above in side-captured manner. The dataset was collected from 23 subjects under three different lighting conditions. All subjects were within a specific age range; they had no disabilities, or any other problems that could affect driving. Each subject performed all 13 motions in a single recording, where each motion was recorded for around 3 seconds under the same brightness condition. The three brightness conditions were classified as "bright," "moderate," and "dark" to test how much brightness would affect the recordings and to examine the robustness of event data in harsh illumination condition. We note that informed consent was obtained from all participants during data collection. To evaluate the classifier’s performance, we randomly selected 992 motions for the training set, and the remaining 247 motions were designated as the test set.
Due to limitations in recruiting participants, we were unable to include experiments involving drivers with medical conditions or disabilities. This will be expanded in future work. Instead, to address unpredictable scenarios, we introduced five new types of event data that were not used during training. These additional test cases include abnormal reactions such as seizures during driving, sudden standing or rising movements, collisions where the driver hits the windshield, hand-swipe gestures, and abnormal camera angles caused by accidents or issues with the camera’s fixed position. These newly created event-based video sequences were specifically designed to test how the system responds to other abnormal cases not seen during training.
4.3 EXPERIMENTAL SETUP AND TOOL FLOW
The experiment is mainly consisted of 3 sub-parts to measure different performance of our proposed model: the accuracy of the 13-class driver motion classification, the accuracy of classifying unexpected abnormal driver motion versus normal driver motion, and the energy efficiency with respect to throughput in various AI accelerators.
4.3.1 Multiple driver motion classification
The aim of this experiment lies in observing how accurately our proposed system can classify various driver motions (13 classes in total). With the model developed based on the Lava-DL API, our proposed network will be loaded onto the RTX 3080 GPU and trained for 200 epochs. During each epoch in training, the models would be evaluated with the test set and produce accuracy based on the spike rates of the output spikes. If the model produces the highest test accuracy, the model’s weights are stored as a PyTorch state dictionary file. Further explanation about the configuration of the system is shown in subsection 4.4.
Figure 5: Sample shots taken of abnormal driver gestures. Note that hand-swipe gestures are not simple hand-waving actions but are actions to protect oneself from various attacks
Table 2: Abnormal driver gestures
Seizures during driving
Sudden standing
Collisions where the driver hits the windshield
Hand-swipe gestures
Abnormal camera angles
4.3.2 Inference on unseen abnormal driving motions
While it is important to see how the model performs classifications based on the provided 13 motion sets, it is also important to check the model’s capability in classifying unexpected motion sets. Here, we make inferences on a separate dataset that is composed of abnormal driver motions not included in the 13 classes, as well as normal motions, where the model has been trained to classify them as label 3. In this experiment, we have set the objective as binary classification: abnormal driver motion and normal driving motion. Given such abnormal driver motions, if the model produces predictions other than label 3 (normal driving motion), we can observe that the model is capable of classifying them as abnormal motion, and vice versa. We have used the fully trained CSNN model with the same configuration for event data (e.g., sampling time) that was used for measuring accuracy on the 13-class classification task. The test data for this task is composed of 82 abnormal driver motions (10 participants performing 5 different unexpected abnormal driving motions, as shown in figure 5 and table 2) along with 63 normal driver motions from the original test set. The evaluation process was conducted in the same way as the multiple driver motion classification, except in a binary classification fashion.
4.3.3 Energy efficiency in various AI accelerators
Energy efficiency is one of the most promising aspects of SNN, and this experiment is designed to show the overall efficiency of the model with respect to 3 AI accelerators: Nvidia RTX 3080, Nvidia Jetson Xavier NX, and Intel Loihi 2. For both Nvidia GPUs, we loaded the models directly onto the devices and measured various inference costs (latency, total energy consumption, etc.). Due to fast inference speed of GPUs, we measured by averaging each of the total costs by the number of samples processed to measure the values. For Loihi 2, we designed a specific pipeline that enables us to load the model into the system and feed the event data directly to the model, along with the utilization of measurement tools provided by the Lava API. Detailed information about experimental flows and settings for Jetson Xavier NX and Loihi 2 is described in subsection 4.5.